OpenStudy (anonymous):

A parallel circuit has a 5 ohm resistor, 10 ohm resistor and a 15 ohm resistor. What is the equivalent resistance? I know you use the formula: 1/R(eq) = 1/R(1) + ... But I get the fraction 11/30. Is the R(eq) 30? Because the numerator is 11._. Help?

OpenStudy (anonymous):

No, 1/R = 11/30. So R = 30/11 ohms.

OpenStudy (anonymous):

Oh, Thank you:D!